OpenStudy (anonymous):

Which is an equation in point-slope form for the given point and slope? Point: (–3, 7); Slope: 4

OpenStudy (anonymous):

y=mx+c

OpenStudy (anonymous):

y=mx+b we know slope so m=4 and we know the x and y coords so x=-3 and y=7 y=mx+b 7=4(-3)+b 7=-12+b 7+12=b b=19 therefore y=4x+19

OpenStudy (anonymous):

A. y – 7 = 4(x – 3) B. y – 7 = 4(x + 3) C. y + 7 = 4(x + 3) D. y + 7 = 4x – 12

OpenStudy (anonymous):

does the answers

OpenStudy (anonymous):

@laynamae

OpenStudy (anonymous):

I'm so sorry :( I have absolutely no clue ... Good luck though :)

OpenStudy (anonymous):

all man

OpenStudy (gorv):

y-7=4(x+3)

OpenStudy (gorv):

y-y1=\[y-y1=\frac{ y2-y1 }{ x2-x1 }*(x-x1)\]

OpenStudy (gorv):

it is point slope form

OpenStudy (gorv):

\[\frac{ y2-y1 }{ x2-x1 } is slope here\]

OpenStudy (gorv):

as we already have slope =4 so no need to go for slope and (x1,y1 ) are (-3,7) @zobobozo